A basketball player has scored 10 free throws out of 18 tries. He really wants to bring his free throw average up to at least 80%. How many consecutive free throws must he score in order to bring up his average to 80%?
I know this has something to do with rational equations, because the whole chapter we have just done for algebra 2 was about rational equations. Including: internationalizing a denominator, complex fractions, using least common denominator to simplify proportions to solve by cross multiplying. How do you solve this problem?

Answer by Theo(13342)

his average is now 10/18.

he wants to bring it up to 80%.

he will toss another x times.

since his average needs to be 80%, and since 80% is the same as .8 which is the same as 8/10, then:

(x + 10) / (x + 18) = 8/10

to solve this euatio0n, do the following:

cross multiply to get:

10 * (x + 10) = 8 * (x + 18)

simplify to get:

10x + 100 = 8x + 144

subtract 8x from both sides of the equation and subtract 100 from both sides of the equation to get:

10x - 8x = 144 - 100

simplify to get:

2x = 44

divide both sides of the equation by 2 to get:

x = 22

that's your solution.

(10 + x) / (18 + x) becomes (10 + 22) / ( 18 + 22) which becomes 32/40.

multiply 32/40 by (1/4) / (1/4) to get 8/10 is the same as .8 which is the same as 80%.